A nucleus has a radius of 4 Fermis. Use the uncertainty principle to estimate the kinetic energy for a neutron localized inside the nucleus. Do the same for an electron. 


1
Expert's answer
2022-03-23T13:59:54-0400

Answer

Uncertainty of kinetic energy is

"=\frac{\hbar}{2\pi R}\\=\frac{6.626*10^{-34}}{2*3.14*4*10^{-15}}\\=1.06*10^{-19}J"